# Multi-type branching processes with immigration generated by point processes

### v1

@@ Abstract
Recently, the case where the times of immigration are generated by a non-homogeneous Poisson process was considered in depth. In this work, we try to demonstrate how one can use the framework of point processes in order to go beyond the Poisson process. As an illustration, we show how to transfer techniques from the case of Poisson immigration to the case where it is spanned by a determinantal point process.
@@ 1 Introduction
The main purpose of this paper is to examine the impact of immigration on behaviour of the continuous-time Markov branching processes, both single- and multi-type, focussing on immigration types that differ from the Poisson model. Although classical Poisson processes are the most commonly used to model the immigration component, they have certain limitations in practical applications. These include the assumption of equidispersion, which may not always hold, and their time-homogeneity, which is addressed by introducing time-non-homogeneous intensity rates. Recently, such branching models with immigration at the jump times of a non-homogeneous Poisson process were considered in depth by Mitov et al. [17] in the critical case and later on by Slavtchova-Bojkova et al. [42, 43] in the non-critical cases. In these studies, the efforts were concentrated mainly on analysing the asymptotic behaviour of the processes for various rates of the Poisson measure, under the assumption that these rates are asymptotically equivalent to either exponential or regularly varying functions. Consequently, results similar to the strong law of large numbers and central limit theorems were established.
Multi-type Markov branching processes were first introduced by Kolmogorov and Dmitriev [20], marking the term branching process as one of the earliest concepts in the literature. The notion of branching processes with immigration was later formalised by Sevastyanov [38], who explored a single-type Markov process with immigration driven by a time-homogeneous Poisson process, and derived limiting distributions for subcritical, supercritical, and critical cases.
@@
Branching processes with time-non-homogeneous immigration were first introduced by Durham [10] and Foster and Williamson [12]. Comprehensive reviews of these models can be found in the monographs by Badalbaev et al. [4] and Rahimov [36], as well as in Rahimov’s more recent review paper [37]. In the recent work of Rabehasaina and Woo [35], the model described in Mitov et al. [17] and Slavtchova-Bojkova et al. [42, 43] were considered and the established limit results were acquired by means of characteristic functions.
In this work, we aim to introduce a general framework for the Laplace transforms of random point processes within the context of branching processes. We believe that this approach will be valuable for researchers looking to extend their work beyond the Poisson process. For example, our method simplifies the derivation of the generating function for a branching process with immigration, which often serves as the foundation for many studies. Traditionally, this derivation was based on the exact distribution of points over a given interval, which is tractable in the case of a Poisson process, as demonstrated in [17, Theorem 1], [35, Lemma 3.1], and [7, Theorem 1]. Our framework provides a more unified and often simpler approach to addressing problems involving immigration, as one can see from the proof of the third item of Theorem 4.5 and similar results for Poisson immigration, such as [17, Theorem 8] and [35, Theorem 3]. Due to the tractability of the framework, we correct minor errors from the last two referenced results, see Remark 4.6.
The paper is structured as follows: in Section 2 we define the framework of point processes, and give a few examples of such objects, and in Section 3 we introduce single- and multi-type branching processes with immigration. Section 4 contains results for the probability generating functions of the branching process with immigration spanned by a general point process, and asymptotic results in the case where the immigration is generated by a Determinantal Point Process (DPP). More precisely, in Theorem 4.1 the generating functions of branching processes with immigration spanned by DPPs, Cox process and Fractional Poisson Process (FPP) are derived in the single-type case, while Theorem 4.2 presents the multivariate analogue. In Subsection 4.3, for the case of DPPs immigration, we establish the equations for the mean and covariances of the process. Furthermore, in Theorem 4.4, we prove limit results analogous to those in the case without immigration, covering sub-, super-, and critical cases, respectively. Section 5 contains the proofs of the new results.
@@ 2 Point processes on the real line
As we will be interested in point processes (PPs) describing the times of immigrants joining a branching process, we define the basic notions for PPs on $\mathbb{R}$ (or just PPs from now on) although the formalism for defining them on an arbitrary Polish space is similar.
@@
We follow the presentation of [22, Chapter 2.1], however, another suitable reference is the excellent monograph by Baccelli et al. [3].
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, $\mathcal{B}(\mathbb{R})$ be the usual Borel $\sigma$-algebra on $\mathbb{R}$, $\mathbf{N}_{<\infty}$ be the set of measures on $\mathbb{R}$ such that for each $B\in\mathcal{B}(\mathbb{R}),\Phi(B)\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$, and $\mathbf{N}$ the set of measures which can be represented as a countable sum of elements of $\mathbf{N}_{<\infty}$. Let $\mathcal{N}$ be the $\sigma$-algebra generated by the sets
$\{\Phi\in\mathbf{N}:\Phi(B)=k\text{ for some }B\in\mathcal{B}\text{ and }k\in\mathbb{N}_{0}\}.$
@@
where $\delta_{x}$ is the Dirac mass at $x$. For a deeper and more general presentation, see [22, Chapter 2.1], [3, Chapter 1] or [18, Chapter 2].
@@
  1. i)
    its intensity measure $\Lambda$, defined by $\Lambda(B)=\mathbb{E}\left[\Phi(B)\right]$, where $B\in\mathcal{B}(\mathbb{R})$;
  2. ii)
@@
    $\mathcal{L}_{\Phi}(f):=\mathbb{E}\left[e^{-\int f\mathrm{d}\Phi}\right]=\mathbb{E}\left[e^{-\sum_{i}f(X_{i})}\right].$ (2.1)
@@
  3. iii)
    its joint intensities $\rho_{k}$, if they exist, defined as $\rho_{k}:\mathbb{R}_{+}^{n}\to\mathbb{R}_{+}$ such that for each disjoint $B_{1},\dots,B_{k}\in\mathcal{B}(\mathbb{R})$,
    $\mathbb{E}\left[\prod_{i=1}^{k}\Phi(B_{i})\right]=\int_{B_{1}\times\dots\times B_{n}}\rho_{k}(x_{1},\dots,x_{k})\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{k}).$
@@ 2.1 Poisson processes
  1. i)
    for every $B\in\mathcal{B}(\mathbb{R})$, $\Phi(B)\sim Pois(\Lambda(B))$ and
  2. ii)
    for every disjoint $B_{1},\dots,B_{m}\in\mathcal{B}(\mathbb{R})$, $\Phi(B_{1}),\dots,\Phi(B_{n})$ are independent.
It is a consequence that the Laplace transform of this PP is then
$\mathcal{L}_{\Phi}(f)=\exp\left(-\int_{\mathbb{R}}\left(1-e^{-f(x)}\right)\Lambda(\mathrm{d}x)\right).$
In the following three sections, we list some PPs which are often used and cover, respectively, repulsive, clustering, and heavy-tailed behaviour of the interarrival times.
@@ 2.2 Determinantal point processes (DPPs)
DPPs were introduced by Macchi [25] under the name fermion process due to their repulsive behaviour. Since then, they appear in various contexts as describing fundamental objects in random matrix theory, random analytic functions, statistical mechanics, and machine learning [21]. For an introduction to their formalism, we refer to [3, Chapter 5], and for an overview, to the survey [8].
We will call a point process (in this work mainly on $\mathbb{R}_{+}$) $(\Lambda,K)$-determinantal with a directing measure $\Lambda$ if its joint densities satisfy, for $x_{i}\in\mathbb{R}$,
$\rho_{k}(x_{1},\dots,x_{k})=\det(K(x_{i},x_{j}))_{1\leq i,j\leq k}$
with $K$ a Hermitian, locally trace class, integral operator on $L^{2}(\mathbb{R},\Lambda)$. Its Laplace transform is given, for a function $f:\mathbb{R}_{+}\to\mathbb{R}$ with bounded support and
$\varphi_{n}(x_{1},\dots,x_{n}):=\prod_{i=1}^{n}(1-e^{-f(x_{i}))}),$
@@
$\mathcal{L}_{\Phi}(f)=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{\mathbb{R}^{n}}\varphi_{n}(x_{1},\dots,x_{n})\rho_{n}(x_{1},\dots,x_{n})\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n}),$ (2.2)
see [3, Proposition 5.1.18, Corollary 5.1.19]. In particular, note that if we choose the kernel $K(x,y)=\mathbbm{1}_{\left\{x=y\right\}}$, then $\rho_{n}\equiv 1$ for all $n$, and we obtain the usual $\Lambda$-Poisson process.
@@ 2.3 Cox Processes
Cox processes, also known as doubly stochastic Poisson processes, are commonly used for modelling phenomena where event occurrence is influenced by underlying random factors, as demonstrated in [31, 32, 48] and more recently in [9], which discuss their applications in spatial and spatio-temporal data analysis. These processes exhibit behaviour that contrasts with the repulsive nature of determinantal point processes (DPPs), making them particularly suitable for capturing patterns where events tend to cluster.
Formally, a Cox process is constructed by first drawing an intensity measure $\Lambda$ from a random distribution $\eta$ defined on the space of $\sigma$-finite measures on $\mathbb{R}$. Given this intensity measure, a Poisson process with intensity $\Lambda$ is then generated. Formally, we can describe this construction as:
$\Lambda$ $\sim\eta,$
[…]
@@ 2.4 Fractional Poisson processes
The fractional Poisson process (FPP) was introduced by Mainardi et al. [26] as a generalisation of the standard homogeneous Poisson process, which is non-Markovian and has heavy-tailed interarrival times; see also the monograph by Meerschaert and Sikorskii [28]. This process is of renewal type, that is, we can consider that the atoms of the associated random measure are situated at $X_{n}=T_{1}+\dots+T_{n}$ with iid $T_{i}$, such that $\mathbb{P}(T_{n}>t)=E_{\beta}(-\lambda t^{\beta})$ for $\beta\in(0,1)$ and
$E_{\beta}(z):=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(1+\beta k)}$
denote the Mittag-Leffler function. It should be noted that $\mathbb{E}\left[T_{i}\right]=\infty$ and that by the work of Meerschaert et al. [27], the counting process $N_{t}:=\Phi((0,t])$ can be constructed as time-changed $\Lambda$-Poisson process, for $t\geq 0$ and $\alpha\in(0,1)$,
$N_{\alpha}(t)=N(Y_{\alpha}(t)),$ (2.3)
where $N$ is the classical homogeneous Poisson process with parameter $\lambda>0$, and $Y_{\alpha}$ is the inverse of an independent $\alpha$-stable subordinator.

### v2

@@ Abstract
Recently, the case where the times of immigration are generated by a non-homogeneous Poisson process was considered in depth. In this work, we demonstrate how one can use the framework of point processes in order to go beyond the Poisson process. As an illustration, we show how to transfer techniques from the case of Poisson immigration to the case where it is spanned by a determinantal point process.
@@ 1 Introduction
The main purpose of this paper is to examine the impact of immigration on the behaviour of both single- and multitype continuous-time Markov branching processes, focussing on immigration types that differ from the Poisson model. Although classical Poisson processes are the most commonly used to model the immigration component, they have certain limitations in practical applications. These include the assumption of equidispersion, which may not always hold, and their time-homogeneity, which is addressed by introducing time-non-homogeneous intensity rates. Recently, such branching models with immigration at the jump times of a non-homogeneous Poisson process were considered in depth by Mitov et al. [17] in the critical case, and later by Slavtchova-Bojkova et al. [44, 45] in the non-critical case. In these studies, the efforts were concentrated mainly on analysing the asymptotic behaviour of the processes for various rates of the Poisson measure, under the assumption that these rates are asymptotically equivalent to either exponential or regularly varying functions. Consequently, results similar to the strong law of large numbers and central limit theorems were established.
Multitype Markov branching processes were first introduced by Kolmogorov and Dmitriev [21], marking the term branching process as one of the earliest concepts in the literature. The notion of branching processes with immigration was later formalised by Sevastyanov [39], who explored a single-type Markov process with immigration driven by a time-homogeneous Poisson process, and derived limiting distributions for subcritical, supercritical, and critical cases.
@@
Branching processes with time-non-homogeneous immigration were first introduced by Durham [10] and Foster and Williamson [12]. Comprehensive reviews of these models can be found in the monographs by Badalbaev et al. [4] and Rahimov [37], as well as in Rahimov’s review paper [38]. In the recent work of Rabehasaina and Woo [36], the model described in Mitov et al. [17] and Slavtchova-Bojkova et al. [44, 45] was considered, and the established limit results were obtained by means of characteristic functions.
In this work, we aim to introduce the general framework of Laplace functionals of random point processes within the context of branching processes. We believe that this approach will be valuable for researchers looking to extend their work beyond the Poisson process. For example, our method simplifies the derivation of the generating function for a branching process with immigration, which often serves as the foundation for many studies. Traditionally, this derivation was based on the exact distribution of points over a given interval, which is tractable in the case of a Poisson process, as demonstrated in [17, Theorem 1], [36, Lemma 3.1], and [7, Theorem 1]. Our framework provides a more unified and often simpler approach to addressing problems involving immigration, as one can see from the proof of the third item of Theorem 5.3 and similar results for Poisson immigration, such as [17, Theorem 8] and [36, Theorem 3]. Due to the tractability of the framework, we correct minor errors from the last two referenced results, see Remark 5.4.
The paper is structured as follows: in Section 2 we define the framework of point processes, and give a few examples of such objects, and in Section 3 we introduce single- and multitype branching processes with immigration. Section 4 contains results for the probability generating functions of the branching process with immigration spanned by a general point process, and asymptotic results in the case where the immigration is generated by a Determinantal Point Process (DPP). More precisely, in Theorem 4.1 the generating functions of branching processes with immigration spanned by DPPs, Cox process and Fractional Poisson Process (FPP) are derived in the single-type case, while Theorem 4.2 presents the multivariate analogue. In Subsection 5.1, for the case of DPP immigration, we establish equations for the mean and covariances of the process. Furthermore, in Theorem 5.2, we prove limit results analogous to those in the case without immigration, covering sub-, super-, and critical cases, respectively. Section 6 contains the proofs of the new results.
@@ 2 Point processes on the real line
As we will be interested in point processes (PPs) describing the times of immigrants joining a branching process, we define the basic notions for PPs on $\mathbb{R}_{+}:=[0,\infty)$ (or just PPs from now on) although the formalism for defining them on an arbitrary Polish space is similar.
@@
We follow the presentation of [23, Chapter 2.1], however, another suitable reference is the excellent monograph by Baccelli et al. [3].
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, $\mathcal{B}(\mathbb{R})$ be the usual Borel $\sigma$-algebra on $\mathbb{R}$, and $\mathbf{N}_{<\infty}$ be the set of measures $\widetilde{\Phi}$ on $\mathbb{R}_{+}$ such that for each $B\in\mathcal{B}(\mathbb{R}_{+}),\widetilde{\Phi}(B)\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. Further, denote by $\mathbf{N}$ the set of measures which can be represented as a countable sum of elements of $\mathbf{N}_{<\infty}$, and let $\mathcal{N}$ be the $\sigma$-algebra generated by the sets
$\{\Phi\in\mathbf{N}:\Phi(B)=k\text{ for some }B\in\mathcal{B}(\mathbb{R}_{+})\text{ and }k\in\mathbb{N}_{0}\}.$
@@
where $\delta_{x}$ is the Dirac mass at $x$, so $\Phi$ places unit mass at the random locations $X_{1},X_{2},\dots,X_{\kappa}$. For a deeper and more general presentation, see [23, Chapter 2.1], [3, Chapter 1] or [18, Chapter 2].
@@
  • its intensity measure $\Lambda$, defined by $\Lambda(B)=\mathbb{E}\left[\Phi(B)\right]$, where $B\in\mathcal{B}(\mathbb{R}_{+})$;
@@
    $\mathcal{L}_{\Phi}(f):=\mathbb{E}\left[e^{-\int f\mathrm{d}\Phi}\right]=\mathbb{E}\left[e^{-\sum_{i\leq\kappa}f(X_{i})}\right].$ (2.1)
@@
  • its joint intensities $\rho_{k}$, if they exist, defined as $\rho_{k}:\mathbb{R}_{+}^{k}\to\mathbb{R}_{+}$ such that for each disjoint $B_{1},\dots,B_{k}\in\mathcal{B}(\mathbb{R}_{+})$,
    $\mathbb{E}\left[\prod_{i=1}^{k}\Phi(B_{i})\right]=\int_{B_{1}\times\dots\times B_{k}}\rho_{k}(x_{1},\dots,x_{k})\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{k}).$
@@ 2.1 Poisson processes
  1. 1.
    for every $B\in\mathcal{B}(\mathbb{R}_{+})$, $\Phi(B)\sim Pois(\Lambda(B))$ and
  2. 2.
    for every disjoint $B_{1},\dots,B_{m}\in\mathcal{B}(\mathbb{R}_{+})$, $\Phi(B_{1}),\dots,\Phi(B_{m})$ are independent.
It is a consequence that the Laplace functional of this PP is then
$\mathcal{L}_{\Phi}(f)=\exp\left(-\int_{\mathbb{R}_{+}}\left(1-e^{-f(x)}\right)\Lambda(\mathrm{d}x)\right).$ (2.2)
If $\Lambda(\mathrm{d}x)=\lambda\mathrm{d}x$, then we say that the Poisson process is homogeneous of rate $\lambda$.
In the following three subsections, we list some PPs which are often used and cover, respectively, repulsive, clustering, and heavy-tailed behaviour of the interarrival times.
@@ 2.2 Determinantal point processes (DPPs)
Determinantal point processes (DPPs) were introduced by Macchi [26] under the name fermion process, due to their repulsive behaviour. Since then they have arisen in various contexts including random matrix theory, zeros of random analytic functions, statistical mechanics, and even machine learning [22]. We refer the reader to [3, Chapter 5] for an introduction of their formalism, and to the survey [8] for an overview.
We call a point process $\Phi$ on $\mathbb{R}_{+}$ $(\Lambda,K)$-determinantal if $\Lambda$ is a locally finite measure on $\mathbb{R}_{+}$, and its joint intensities satisfy, for $\boldsymbol{x}=(x_{1},\dots,x_{n})$,
$\rho_{n}(x_{1},\dots,x_{n})=\det\left(K(x_{i},x_{j})\right)_{1\leq i,j\leq n}=:D(\boldsymbol{x}),$
where $K\colon\mathbb{R}_{+}^{2}\to\mathbb{R}_{+}$ is symmetric; for $\Lambda^{n}$-almost all $\boldsymbol{x}$, $\left(K(x_{i},x_{j})\right)_{1\leq i,j\leq n}$ is non-negative definite, and for each bounded $D\in\mathcal{B}(\mathbb{R}_{+})$, $\int_{D}K(x,x)\Lambda(\mathrm{d}x)$ is finite. These properties ensure that $\Lambda$ and $K$ determine uniquely $\Phi$, see [3, Corollary 5.1.14]. However, they may look too implicit, as they do not describe how to build admissible kernels. One recipe is to look at regular kernels of the type $K(x,y)=\sum_{n\in\mathbb{N}}\lambda_{n}\phi_{n}(x)\phi_{n}(y)$ for $\phi_{i}$ orthonormal in $L^{2}(\Lambda,\mathbb{R}_{+})$ and $\lambda_{i}\in[0,1]$, see [3, Theorem 5.2.5].
Further, the Laplace functional of $\Phi$ is given, for any nonnegative $f$ of compact support, and
$\varphi_{n}(\boldsymbol{x}):=\prod_{i=1}^{n}(1-e^{-f(x_{i}))}),\quad\quad\Lambda(\mathrm{d}\boldsymbol{x}):=\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n})$
@@
$\begin{split}\mathcal{L}_{\Phi}(f)&=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{\mathbb{R}_{+}^{n}}\rho_{n}(x_{1},\dots,x_{n})\varphi_{n}(x_{1},\dots,x_{n})\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n})\\ &=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{\mathbb{R}_{+}^{n}}D(\boldsymbol{x})\varphi_{n}(\boldsymbol{x})\Lambda(\mathrm{d}\boldsymbol{x}),\end{split}$ (2.3)
see for example [3, Proposition 5.1.18 and Corollary 5.1.19] or [42, Theorem 3.6 with $\alpha=-1$]. In particular, note that if $\Lambda(\mathrm{d}x)$ is diffuse and we choose the kernel $K(x,y)=\mathbbm{1}_{\left\{x=y\right\}}$, then for all $n$, $\rho_{n}$ is equal to 1 $\Lambda(\mathrm{d}\boldsymbol{x})$-almost everywhere, so we obtain the usual $\Lambda$-Poisson process, see [3, Example 5.1.6] for a rigorous derivation.
@@ 2.3 Cox Processes
Cox processes, also known as doubly stochastic Poisson processes, extend the ordinary Poisson process by randomising its intensity measure. They are usually used for modelling phenomena where event occurrence is influenced by underlying random factors, as demonstrated in [32, 33, 50] and more recently in [9], which discuss their applications in spatial and spatio-temporal data analysis.
Formally, let $\eta$ be a random $\sigma$-finite measure on $\mathbb{R}_{+}$. Then we call a point process $\Phi$ a Cox process with directing measure $\eta$ if conditional on $\eta$
$\Phi\mid\eta\sim\text{$\eta$-Poisson process}.$
Therefore, its Laplace functional is obtained by averaging over $\eta$ in (2.2)
$\mathcal{L}_{\Phi}(f)=\mathbb{E}\left[\exp\Bigl(-\int_{\mathbb{R}_{+}}\bigl(1-e^{-f(x)}\bigr)\,\eta(\mathrm{d}x)\Bigr)\right],$ (2.4)
for every nonnegative measurable $f$. The first two moment measures follow by conditioning on $\eta$ and using the Poisson moment formula, as in [23, Proposition 13.6]:
$\mathbb{E}\left[\Phi(B)\right]=\mathbb{E}\left[\eta(B)\right],\qquad\operatorname{Var}\bigl(\Phi(B)\bigr)=\operatorname{Var}\bigl(\eta(B)\bigr)\;+\;\mathbb{E}\left[\eta(B)\right],$
for all $B\in\mathcal{B}(\mathbb{R}_{+})$. For proofs and more properties, see for example [23, Chapter 13].
@@ 2.4 Fractional Poisson processes
The fractional Poisson process (FPP), introduced by Mainardi et al. [27] is a non-Markovian generalisation of the standard homogeneous Poisson process, which has heavy-tailed interarrival times; see also the monograph by Meerschaert and Sikorskii [29]. It depends on parameters $\beta\in(0,1]$ and $\lambda>0$, and is of renewal type, that is, the atoms of the associated random measure $\Phi_{\beta,\lambda}$ are situated at $X_{n}=T_{1}+\dots+T_{n}$ with iid $T_{i}$ such that
$\mathbb{P}(T_{i}>t)=E_{\beta}(-\lambda t^{\beta}),\quad\text{where}\quad E_{\beta}(z):=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(1+\beta k)}$
is the Mittag-Leffler function. Define the counting process by $N_{\beta,\lambda}(t):=\Phi_{\beta,\lambda}((0,t])$. For $\beta=1$, we get $T_{i}\sim Exp(\lambda)$, so the FPP is in fact a homogeneous Poisson process of rate $\lambda$. However, for $\beta\in(0,1)$, $\mathbb{P}(T_{i}>t)$ is of order $1/t^{\beta}$ and a fortiori the expectation $\mathbb{E}\left[T_{i}\right]$ is infinite. The work of Meerschaert et al. [28] extends the link to Poisson processes beyond $\beta=1$ by showing that
$N_{\beta,\lambda}(t)=\mathcal{N}_{\lambda}(Y_{\beta}(t)),$ (2.5)
where $\mathcal{N}_{\lambda}$ is a homogenous Poisson process of rate $\lambda>0$, and $Y_{\beta}$ is the inverse of an independent $\beta$-stable subordinator. Thus understanding the asymptotic properties of inverse subordinators $Y_{\beta}$, for example as in [1], is key to deriving limit theorems and other results for the FPP. Consequently, the Laplace functional of $\Phi_{\beta,\lambda}$ can be written
$\mathcal{L}_{\Phi_{\beta,\lambda}}(f)=\mathbb{E}\left[\exp\left(-\lambda\int_{\mathbb{R}_{+}}\left(1-e^{-f(t)}\right)\mathrm{d}Y_{\beta}(t)\right)\right].$ (2.6)